Ten artists (of two different experience levels) are exhibiting their paintings in an art gallery. Three have exhibited in previous shows (making them “experienced artists”) and seven have not. If a patron selects three paintings to purchase at random, find the probability of selecting:

(A) the probability of selecting at least 2 experienced artists’ paintings.

(B) the probability of selecting no experienced artists’ paintings.

(C) the probability of selecting some experienced artists’ paintings.

(D) the probability of selecting less than 2 experienced artists’ paintings.
Answer #1

P (selecting experienced artists painting) = B = + (say) patron selects 3 paintings Now a random. let Х be selected randomc) Probability of selecting some experienced artists paintings = P(x²1) = 1- P(x=0) 1- 0.343 = 0.65 7. ress than 2 Probabili